OpenStudy (anonymous):

What is the sum of three and three over four plus two and three over four in simplest form?

OpenStudy (mysangels):

What do you mean by "and"? Is it 3+3 + 2+3 4 4

OpenStudy (anonymous):

what is the sum of 3 3/4 plus 2 3/4 in simplest form pls

OpenStudy (anonymous):

I'll do it @jennawenna13

OpenStudy (anonymous):

you can do it i just need help

OpenStudy (anonymous):

= 15/4 + 11/4 = ((15 × 4) + (11 × 4)) / (4 × 4) = (60 + 44) / 16 = 104/16 = 13/2 ANSWER: 3 3/4 + 2 3/4 = 6 1/2

OpenStudy (anonymous):

thank u vary much :)

OpenStudy (anonymous):

no problem hun just send me a message or tag me if you have any other questions :)

OpenStudy (anonymous):

and also make sure you close the question in case you have any other questions to post